An unknown material that is supposedly good for wiring has been brought to your attention. If 100m of .01mm diameter wire has a resistivity of 5 ohms, what is its resistivity?
Answer #1

A = pi*r^2

r = 1/2 * 0.01*10^-3 m

R = rho *L /A

5 = rho * 100 / ( pi * 0.5 * 0.01 *10^-3 )^2

solve for rho = 1.2337*10^-11   ohmmetre